OpenStudy (anonymous):

Multiply ans simplify. What expression is in front of the radical in your simplified answer. 2√x* 3√y *2√z^2

OpenStudy (mertsj):

\[2\sqrt{x}(3\sqrt{y})(2\sqrt{z^2)}=2(3)(2)\sqrt{x}\sqrt{y}\sqrt{z^2}\]

OpenStudy (mertsj):

Perhaps you can finish it.

OpenStudy (anonymous):

sqrt{z^2}=z

OpenStudy (anonymous):

I don;t understrand all of the { [

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

The parentheses ( )( ) means to multiply

OpenStudy (mertsj):

The others are radicals which means "take the second root"